OpenStudy (anonymous):

PLEASE HELP I NEED TO GET THIS ASSIGNMENT DONE! A company produces accessories for smart phones and tablets. The profit on each smart phone case is $2, and the profit on each tablet case is $3. The company made a profit of $1,200 on the cases last month. The equation 2x + 3y = 1,200 represents the company's profit from cases last month, where x is the number of smart phone cases sold and y is the number of tablet cases sold.

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

2.Describe how you would graph this line using the slope-intercept method. Be sure to write in complete sentences.

OpenStudy (anonymous):

3.Write the equation in function notation. Explain what the graph of the function represents. Be sure to use complete sentences

OpenStudy (anonymous):

4.Graph the function using one of the following two options below. On the graph, make sure to label the intercepts. You may graph your equation by hand on a piece of paper and scan your work. You may graph your equation using graphic technology that can be found in the Course Information area.

OpenStudy (anonymous):

4.Suppose in the next month, the total profit on smart phone cases and tablet cases is $1,500. The profit amounts are the same, $2 for smart phone case and $3 for the tablet case. In a paragraph of at least three sentences, explain how the graphs of the functions for the two months are the same and how they are different. Be sure to use complete sentences.

OpenStudy (texaschic101):

1. slope intercept form is y = mx + b 2x + 3y = 1200 --- subtract 2x from both sides 3y = -2x + 1200 --- divide both sides by 3 y = -2/3x + 400 In y = mx + b form, the slope is in the m position and the y intercept is in the b position. So we know the slope is -2/3 and the y intercept is (0,400). To find the x intercept, sub in 0 for y and solve for x y = -2/3x + 400 0 = -2/3x + 400 2/3x = 400 x = 400/(2/3) x = 400 * 3/2 x = 1200/2 which equals 600....so your x intercept is (600,0) so we have x intercept (600,0) and y intercept (0,400). Plot those points. Now start at y intercept , and since slope is -2/3....go down 2 spaces (because it is -2) and to the right 3 spaces, then down 2, and to the right 3...and you should cross the x axis at (600,0).

OpenStudy (texaschic101):

function notation for equation : 2x + 3y = 1200. put it in slope intercept form y = -2/3x + 400 now sub in f(x) in for y f(x) = -2/3x + 400 <== function notation

OpenStudy (anonymous):

@texaschic101 Thank you! my computer was froze